the little girl drank the first part of her bottle in 3 minutes. she drank the second part in 5 minutes. she drank the last part in 2 minutes. what fraction of an hour did the little girl take to drink her bottle of milk?
Answer by stanbon(75887)

the little girl drank the first part of her bottle in 3 minutes. she drank the second part in 5 minutes. she drank the last part in 2 minutes. what fraction of an hour did the little girl take to drink her bottle of milk?
----
time = (3+5+2) = 10 minutes = 6/60 hr = 1/10 hour
-------------------
Cheers,
Stan H.
------------